There is a convex lens whose focal length is 10 cm and the refractive index is 2. If it's one edge is silvered then find the focal length of system

Grade:10

1 Answers

Vikas TU
4 years ago
Dear student 
Here 
1/F = 2/f + 1/fm 
fm = infinity Because it is plano convex lens 
F = 30 
f = 60 after calculation
fm = R/2 
R = 30 cm 
1/f = (mew-1)(1/R)
mew = 1.5 
Good Luck